Two high-current transmission lines carry currents of 28.0A and 80.0A in the same direction and are suspended parallel to each other 38.0cm apart.If the vertical posts supporting these wires divide the lines into straight 16.0

m segments, what magnetic force does each segment exert on the other?
Answer #1

media%2Ff3d%2Ff3d912dc-36f7-479e-9dd4-01

F/L = u * 28*80 / 2* pi * 0.38

F/L = 1.1789*10^-3

F = 16 * 1.1789*10^-3 = 0.0188 N